The length of a rectangle is more than twice its width, and the area of the rectangle is 77 square meters. Find the dimensions of the rectangle.
Answer by RAY100(1637)

Area + Length * Width = w* (>2w) = 77
.
2w^2 <77
.
w^2 < 77/2 < 38.5
.
w^2 < 38.5
.
w< 6.2 m
.
L > 2w,,, L> 12.4
.
check ,,,,Area = L*W = 12.4 * 6.2 = 77,,,,ok
.
Reviewing problem,,,Let w = 6.2,,,,then L > 12.4,,,,,,satisfies requirements